Math 120 Spring 98

Quiz 2

Name:

Basic Calculus 2

Date:  Friday January 30
Time Begun:  Ron Buckmire

Time Ended:

Topic covered: Evaluating the definite integral

Instructions:

1. Once you open the quiz, you have 60 minutes to complete it.

2. You may use the book or any of your class notes, and you may use a calculator. You must work alone.

3. If you use your own paper, please staple it to the quiz before coming to class. If you don’t have a stapler, buy one.

4. After completing the quiz, sign the pledge below stating on your honor that you have adhered to these rules.

5. Relax and enjoy...

6. This quiz is due on Monday, February 2, in class. NO LATE QUIZZES WILL BE ACCEPTED.

Pledge: I, , pledge my honor as a human being and Occidental student, that I have followed all the rules above to the letter and in spirit.

Math 120 Spring 1998 Quiz 2

SHOW ALL YOUR WORK

1. Given

f(x) = \{ \begin{matrix}−1, & if & −4 ≤ x < 0 \\ x − 1, & if & 0 ≤ x ≤ 4 \end{matrix}

(a) (3 points) Sketch the function f(x) on the axes below

(b) (2 points) Use your graph to help you evaluate ∫_{−4}^{0}f(x) dx exactly.

(c) (2 points) Use your graph to help you evaluate ∫_{0}^{4}f(x) dx exactly.

(d) (2 points) Use your previous answers to help you evaluate ∫_{−4}^{4}f(x) dx exactly.

2. (1 point) Now suppose f(x) was defined as f(x) = e^{x^{2}}, −4 ≤ x ≤ 4 instead. Why wouldn’t you be able to use the same technique you used in question (1) to evaluate ∫_{−4}^{4} e^{x^{2}} dx exactly?